OpenStudy (anonymous):

how to solve square root of (a+9)^2

OpenStudy (allank):

Hint: \[(\sqrt{a})^2 = a\]

OpenStudy (anonymous):

a+9?

OpenStudy (allank):

yep

OpenStudy (anonymous):

This symbol means square root - @. What about (2@7-6@5)(3@7+9@5)

OpenStudy (anonymous):

\[(2\sqrt{7}-6\sqrt{5})(3\sqrt{7}+9\sqrt{5})\]

OpenStudy (allank):

Thank you for rewriting that...checking..

OpenStudy (allank):

That will become \[2\sqrt{7}(3\sqrt{7}+9\sqrt{5})-6\sqrt{5}(3\sqrt{7}+9\sqrt{5})\]

OpenStudy (allank):

Then do the multiplication... \[42+18\sqrt{35}-18\sqrt{35}-270\]

OpenStudy (allank):

Are you okay till this point?

OpenStudy (allank):

Still there?

OpenStudy (anonymous):

I am still here. My computer is very slow beacuse it is raining very bad outside. I think they answer to this question if -228?

OpenStudy (allank):

Very correct. Glad you understand. :)

OpenStudy (anonymous):

A little rusty!!! I haven't done algebra in 11 years

OpenStudy (allank):

Wow. that's a long time :)

OpenStudy (anonymous):

Here is another one. t^2-81/(t+9)^2

OpenStudy (anonymous):

Is the answer 0

OpenStudy (allank):

working...

OpenStudy (allank):

sorry, no. it's (t-9)/(t+9)

OpenStudy (allank):

can u identify where you went wrong?

OpenStudy (anonymous):

Yes. I am now working on another problem What is a discrimiant?

OpenStudy (anonymous):

-6x^2+4x-9=0

OpenStudy (allank):

Helping out sum1 else, wait a bit @LilJuneau13